Productivity can be measured in various ways, i.e., by Labor, capital, energy, use of materials, etc. In Modern Lumber Inc. Art Binley, president of this company producing wooden boxes for apples that sells to farmers, has been able, with its current equipment, to produce 240 boxes for every 100 trunks used.  Currently, buy 100 trunks per day and each trunk requires 3 hours of labor to process. Binley believes that you can hire a professional buyer who can purchase high-quality trunks for the same cost. In that case, you can increase your production to 260 boxes per 100 trunks.  Your labor hours increase to 8 per day.

What will be the impact on productivity (measured in boxes per hour of work) if you hire the buyer?
Answer #1

Given:

Number of trunks T = 100

Hours required per trunk H1 = 3 hours

Total hours required = T x H1 = 3 x 100 = 300 hours

Boxes produced = 240

Current productivity P1 = Boxes produced / total hours = 240 / 300 = 0.8 boxes per hour of work

Increase in labor hours = 8 hours

New Total hours required = 300 + 8 = 308 hours

Boxes produced = 260

New productivity P2 = Boxes produced / total hours = 260 / 308 = 0.844 boxes per hour of work

Hence, as we see, the productivity has increased. Increase in productivity = (P2 - P1) *100 / P1 = (0.844-0.8)*100 /0.8 = 0.055 = 5.5% increase = 0.044 boxes per hour of work